OpenStudy (anonymous):

How to smplify (2a-b)^3/(b-2a)^2 I know this one is about law of exponent but the base are not the same so how?

OpenStudy (zzr0ck3r):

note that \(\frac{1}{b-2a}=\frac{-1}{2a-b}\)

OpenStudy (zzr0ck3r):

so we have \(\frac{(2a-b)^3}{(b-2a)}=\frac{-(2a-b)^3}{(2a-b)}=?\)

OpenStudy (zzr0ck3r):

@derick062000 do you understand?

OpenStudy (anonymous):

no i still don't get it, what law did you use? sorry for the late reply i switch laptop thats why it took me loong